OpenStudy (anonymous):

a 3% reduction, how would i put this into a fraction? it was $5.15 now$5.00

OpenStudy (stamp):

100% - 3% = 97% So 97% of $5.15 = $5.00 "%" is "per cent" which is "per 100" or "/100" 97% = 97 per cent = 96 per 100 = 97/100 Substitute the fraction form of 97%: 97/100 (5.15) = 5.00 97/100 (515/100) = 5